OpenStudy (abbycross167):

Can someone please show me the steps to this question? Please You can stuff 108 envelopes in 45 minutes. At this rate, how many envelopes can you stuff in 2 hours?

OpenStudy (kdr1308):

First off, can you find the rate at which how many envelopes are stuffed in 1 minute?

OpenStudy (kdr1308):

x = amount of envelopes per minute t = total envelopes stuffed in two hours 60 in a hour, so 120 in two 108/45 = x --> \[x \times 120 = t\] or \[2(x \times 60) = t\]